OpenStudy (anonymous):

Multiply. (x – 4)(x2 – 3x + 5)

OpenStudy (mertsj):

Multiply each term of the trinomial by x and then multiply each term of the trinomial by -4. Then add the like terms and you are done!!

OpenStudy (anonymous):

6 multiplication steps here, beginning with x times x2 = x^3 x times -3x = -3x^2 x times 5 = 5x Can you get the other three using -4?

OpenStudy (anonymous):

Use foil x(x^(2)) - 3x(x) + 5(x) + (-4)(x^2) + (-3x)(-4) + 5(-4)

OpenStudy (anonymous):

-4 times x2 = ? -4 times -3x = ?? -4 times 5 = ?

OpenStudy (anonymous):

foil with 6 terms?

OpenStudy (anonymous):

x^3-7x^2+17x+20

OpenStudy (mertsj):

-4 times x^2 = -4x^2 -4 times -3x = 12x -4 times 5 = -20

OpenStudy (anonymous):

Can you please start using brackets and this symbol: ^ to denote exponents. Ensuring the exponents are always in brackets. It makes it easier for us to read and understand what you are asking